OpenStudy (anonymous):

Louise buys milk in 2 litre cartons. She uses 1 2/3 (or 5/3) litres of milk every day. What is the least number of cartons of milk she would need to buy in ten days?

OpenStudy (anonymous):

5/3x10=50/3= about 16 liters 9 cartons would be the answer it equals about 18 liters

OpenStudy (anonymous):

Thanks for the help! Where is the 9 cartons from? And how did you get 18 litres?

OpenStudy (anonymous):

She uses 5/3 liters a day. Louise buys 2 liter milk cartons which are equivalent to 6/3. You need to find how much milk she would consume in ten days which would be 50/3 liters which is over 16 liters.

OpenStudy (anonymous):

The least number of cartons she would need to buy is 18 because 8 cartons of milk would be insufficient for the 10 day span.

OpenStudy (anonymous):

How are 2 litre milk cartons equivalent to 6/3? Why is 8 cartons of milk insufficient for the 10 day span? I get the rest of your explanation though, thank you SO SO much! :)

OpenStudy (anonymous):

lol reree 6/3=2

OpenStudy (anonymous):

LOL I feel so stupid!! I was looking at it from a different perspective. Thanks again!